# DiffPatch: Generating Customizable Adversarial Patches using Diffusion Models

### v3

@@ Abstract
Physical adversarial patches printed on clothing can enable individuals to evade person detectors, but most existing methods prioritize attack effectiveness over stealthiness, resulting in aesthetically unpleasing patches. While generative adversarial networks and diffusion models can produce more natural-looking patches, they often fail to balance stealthiness with attack effectiveness and lack flexibility for user customization. To address these limitations, we propose DiffPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches. Our approach allows users to start from a reference image (rather than random noise) and incorporates masks to create patches of various shapes, not limited to squares. To preserve the original semantics during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO). Our method achieves attack performance comparable to state-of-the-art non-naturalistic patches while maintaining a natural appearance. Using DiffPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios. AdvT-shirt-1K can serve as a useful dataset for training or testing future defense methods.
@@ 1 Introduction
Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of DiffPatch in various scenarios.
Object detectors powered by deep neural networks (DNNs) are integral to modern computer vision systems, enabling critical applications such as autonomous driving [9] and medical diagnostics [39]. However, DNNs are highly susceptible to adversarial attacks [8, 57, 16, 55], which manipulate their outputs in ways that defy logic or compromise safety. These attacks can cause object detectors to misclassify or fail to detect critical objects. This vulnerability poses significant risks, particularly in safety-critical domains.
@@
To address these limitations, we propose DiffPatch, a novel adversarial patch generation framework based on Stable Diffusion [45]. DiffPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary. Our approach first maps the reference image into the latent space and optimizes noise latent vectors to create adversarial patches. We employ Null-text inversion [40] to ensure near-perfect reconstruction, preserving the original semantics. Additionally, we introduce an Incomplete Diffusion Optimization (IDO) strategy to balance natural appearance and attack performance, along with an IoU-Detection loss to accelerate convergence. To generate irregular-shaped patches, we use masks to replace the background and suppress gradients in non-essential areas, reducing interference during optimization.
Using DiffPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios. This dataset demonstrates the practicality of our approach and provides a valuable resource for future research.
@@
  • We propose a novel diffusion-based method called DiffPatch for generating naturalistic adversarial patches based on a reference image. DiffPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
  • Experimental results demonstrate that our DiffPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors. Its performance rivals or even surpasses state-of-the-art methods for generating unnatural adversarial patches.
@@ 2.2 Naturalistic Adversarial Patch
Figure 2: An overview of the proposed DiffPatch generation framework.
@@ 3 Proposed Approach
Figure 2 illustrates the overall framework of DiffPatch. It optimizes Null-text embeddings along incomplete diffusion trajectories and maps the image to latent vectors at intermediate timestamps, ensuring that the patch retains a natural appearance during updates. It employs Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, achieving faster convergence and improved attack performance. To address memory constraints, we skip the U-Net gradients during optimization and use target masks to eliminate semantic interference from irrelevant regions, enabling the generation of more naturalistic adversarial patches. Next, we will introduce the key techniques of DiffPatch.
@@ 3.1 Image-to-Latent Mapping
where $w$ is the guidance scale parameter, and $\phi=\mathcal{E}("")$ is the embedding of a null text. The generated content is highly sensitive to textual prompts, and even minor variations can cause trajectory deviations during guided generation. These deviations accumulate, making reconstruction challenging, especially during image editing.
@@ 3.3 Target Mask Control
Overall, the DiffPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance. Finally, Target Mask Control is applied to preserve the natural appearance of the generated patches. The complete generation process is outlined in Algorithm 1. Furthermore, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as illustrated in Figure 3. Specifically, the generated patch can serve as the input for generating a new patch, enabling stronger attack performance while maintaining a natural appearance. This iterative approach ensures both effectiveness and stealthiness in adversarial patch generation.
@@
Algorithm 1 DiffPatch Generation Process
@@
20:  $z_{p}\leftarrow G(\bar{z}_{\frac{T}{2}}+\delta_{B},\frac{T}{2},\mathcal{C},\{\phi_{t}\}^{\frac{T}{2}}_{t=1})\odot m$
@@ 4 Experiments
Table 1: Comparing DiffPatch with other methods in ASR (%) (higher is better). The best results are underlined. denotes transfer attack methods, while the rest are white-box attacks. trained on YLv2, trained on YLv4t. The generated patches are shown at the bottom.
Method ASR ($\uparrow$)
@@
${}^{(P_{A})}$dvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
${}^{(P_{B})}$-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
${}^{(P_{C})}$ 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
${}^{(P_{D})}$2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
${}^{(P_{E})}$ 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
${}^{(P_{F})}$dvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
${}^{(P_{1})}$v3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
${}^{(P_{2})}$v3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
${}^{(P_{3})}$v4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
${}^{(P_{4})}$v4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
${}^{(P_{5})}$v5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
${}^{(P_{6})}$v7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
${}^{(P_{7})}$v10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
${}^{(P_{8})}$v12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
${}^{(P_{9})}$(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
${}^{(P_{10})}$(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
${}^{(P_{11})}$-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
${}^{(P_{j})}$ource 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
@@ 4.1 Experimental Setup
We treat the person as the target object class, setting the batch size to 32, the maximum number of epochs $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, learning rates $\eta_{u}=0.01$ and $\eta_{p}=0.003$, and perturbation constraint $\epsilon=0.5$. We use the Adam optimizer [29] with its default parameters. Textual prompts $\mathcal{P}$ are automatically generated using BLIP [33]. The adopted version of Stable Diffusion [45] is v1.4. Since we generate irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$. To achieve stronger performance, we apply two iterations of optimization to the patch using DiffPatch. We evaluate the effectiveness of different patch generation methods using both the Attack Success Rate (ASR) and Average Precision (AP). The ASR measures the success of the patches in misleading the detector, while AP reflects the overall decline in detection accuracy. The confidence threshold and IoU threshold for the detector are both set to 0.5. All experiments were conducted on a single NVIDIA A100 GPU.
@@ 4.2 Attack Performance Evaluation
We generate a series of adversarial patches targeting different models using DiffPatch and compare them on the INRIA dataset with existing methods, including AdvYL [50], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34]. The detailed results are reported in Table 1. As shown, DiffPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny. When compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YOLOv4 results in a 0.2% and 4.1% ASR increase, respectively. Additionally, as demonstrated in the Appendix B, DiffPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
@@ 4.3 Cross-dataset Evaluation
We also evaluate an ensemble version of DiffPatch, which trains the patch using averaged gradients from multiple models. We compare this with two other methods, NPAP and D2D, where NPAP was trained on YOLOv3 and YOLOv4, and D2D was trained on all considered models. Our ensemble DiffPatch was similarly trained on all models. The results, presented in the Appendix C, demonstrate that our method outperforms NPAP and D2D by a significant margin.
@@ 4.4 Ablation Studies
To ensure the natural appearance of the generated patches, we constrain each optimization step using the $L_{\infty}$ norm. The choice of $\epsilon$ will influence visual naturalness of the generated patch and its attack performance. To quantify the natural appearance of the patches, we use GPT-4o [1] to generate descriptive text for reference images (denoted by ’-’) and compute the CLIP similarity [41] between these descriptions and the generated patches. The corresponding descriptions can be found in the Appendix D. As illustrated in Figure 4, we obtain patches with varying styles under different constraints using DiffPatch and evaluate both the ASR and CLIP similarity for each. Based on the results, we select $\epsilon=0.5$ as the optimal value, as it achieves a favorable balance between attack performance and natural appearance.
@@
Here, we investigate the impact of patch shapes, as used for $P_{m}$ and $P_{n}$ in Table 1. We compare the performance of random noise images with square and irregular shapes (of the same size as DiffPatch). From the ASR and AP results in Table 1, it is evident that square patches yield higher attack effectiveness compared to irregular patches. In other words, designing irregular patches with high attack success rates is more challenging.
@@ 4.5 Adversarial T-shirt Dataset
In total, we collected 1,131 images, including both individual and group photos (ranging from 2 to 10+ persons). We name the dataset AdvT-shirt-1K, which includes detailed annotations for person and patch locations (bounding boxes). The collection and annotation process of the dataset spanned six months, which involves the refinement of our DiffPatch method for multiple times. The statistics of the dataset is shown in Figure 8 with several example images are provided in Figure 1. The results shown in the right subfigure of Figure 8 demonstrate the effectiveness of the adversarial T-shirts in evading the YOLOv5s detector. Although small in scale, our dataset provides a practical resource for developing and testing the real-world effectiveness of adversarial patch detectors.
@@ 5 Conclusion
In this work, we proposed a novel diffusion-based adversarial patch generation method DiffPatch to generate naturalistic and customized patches based on a reference image. DiffPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches. Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch. With DiffPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of DiffPatch in physical-world environments.

### v5

# BadPatch: Diffusion-Based Generation of Physical Adversarial Patches

@@ Abstract
Physical adversarial patches printed on clothing can enable individuals to evade person detectors, but most existing methods prioritize attack effectiveness over stealthiness, resulting in aesthetically unpleasing patches. While generative adversarial networks and diffusion models can produce more natural-looking patches, they often fail to balance stealthiness with attack effectiveness and lack flexibility for user customization. To address these limitations, we propose BadPatch, a novel diffusion-based framework for generating customizable and naturalistic adversarial patches. Our approach allows users to start from a reference image (rather than random noise) and incorporates masks to create patches of various shapes, not limited to squares. To preserve the original semantics during the diffusion process, we employ Null-text inversion to map random noise samples to a single input image and generate patches through Incomplete Diffusion Optimization (IDO). Our method achieves attack performance comparable to state-of-the-art non-naturalistic patches while maintaining a natural appearance. Using BadPatch, we construct AdvT-shirt-1K, the first physical adversarial T-shirt dataset comprising over a thousand images captured in diverse scenarios. AdvT-shirt-1K can serve as a useful dataset for training or testing future defense methods. The code and datasets are available at https://github.com/Wwangb/BadPatch.
@@ 1 Introduction
Figure 1: Example images from our AdvT-shirt dataset, showcasing the performance of BadPatch in various scenarios.
Object detectors powered by deep neural networks (DNNs) are integral to modern computer vision systems, enabling critical applications such as autonomous driving [9] and medical diagnostics [40]. However, DNNs are highly susceptible to adversarial attacks [8, 59, 16, 57, 38, 55], which manipulate their outputs in ways that defy logic or compromise safety. These attacks can cause object detectors to misclassify or fail to detect critical objects. This vulnerability poses significant risks, particularly in safety-critical domains.
@@
To address these limitations, we propose BadPatch, a novel adversarial patch generation framework based on Stable Diffusion [46]. BadPatch generates naturalistic and customizable adversarial patches from a reference image specified by the adversary. Our approach first maps the reference image into the latent space and optimizes noise latent vectors to create adversarial patches. We employ Null-text inversion [41] to ensure near-perfect reconstruction, preserving the original semantics. Additionally, we introduce an Incomplete Diffusion Optimization (IDO) strategy to balance natural appearance and attack performance, along with an IoU-Detection loss to accelerate convergence. To generate irregular-shaped patches, we use masks to replace the background and suppress gradients in non-essential areas, reducing interference during optimization.
Using BadPatch, we construct AdvT-shirt-1K, a dataset of over a thousand images featuring customizable adversarial patches in diverse physical-world scenarios. This dataset demonstrates the practicality of our approach and provides a valuable resource for future research.
@@
  • We propose a novel diffusion-based method called BadPatch for generating naturalistic adversarial patches based on a reference image. BadPatch integrates Null-text inversion, Incomplete Diffusion Optimization (IDO) , masked customization, and an IoU-Detection loss to generate stylized and customizable adversarial patches.
  • Experimental results demonstrate that our BadPatch achieves the highest attack success rate (ASR), significantly reducing the mean average precision (mAP) of various object detectors. Its performance rivals or even surpasses state-of-the-art methods for generating unnatural adversarial patches.
@@ 2.2 Naturalistic Adversarial Patch
Figure 2: An overview of the proposed BadPatch generation framework.
@@ 3 Proposed Approach
Figure 2 illustrates the overall framework of BadPatch. It optimizes Null-text embeddings along incomplete diffusion trajectories and maps the image to latent vectors at intermediate timestamps, ensuring that the patch retains a natural appearance during updates. It employs Incomplete Diffusion Optimization (IDO) guided by an IoU-Detection Loss to optimize the latent vectors, achieving faster convergence and improved attack performance. To address memory constraints, we skip the U-Net gradients during optimization and use target masks to eliminate semantic interference from irrelevant regions, enabling the generation of more naturalistic adversarial patches. Next, we will introduce the key techniques of BadPatch.
@@ 3.1 Image-to-Latent Mapping
[…]
@@ 3.3 Target Mask Control
Overall, the BadPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance. Finally, Target Mask Control is applied to preserve the natural appearance of the generated patches. The complete generation process is outlined in Algorithm 1. Furthermore, the patch can be iteratively optimized using BadPatch to achieve an accumulative effect, as illustrated in Figure 3. Specifically, the generated patch can serve as the input for generating a new patch, enabling stronger attack performance while maintaining a natural appearance. This iterative approach ensures both effectiveness and stealthiness in adversarial patch generation.
@@
Algorithm 1 BadPatch Generation Process
@@
[…]
@@ 4 Experiments
Table 1: Comparing BadPatch with other methods in ASR (%) (higher is better). The best results are underlined. denotes transfer attack methods, while the rest are white-box attacks. trained on YLv2, trained on YLv4t. The generated patches are shown at the bottom.
Method ASR (% $\uparrow$)
@@
$^{(P_{A})}$AdvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
$^{(P_{B})}$T-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
$^{(P_{C})}$NPAP 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
$^{(P_{D})}$D2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
$^{(P_{E})}$DAP 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
$^{(P_{F})}$AdvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
$^{(P_{1})}$YLv3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
$^{(P_{2})}$YLv3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
$^{(P_{3})}$YLv4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
$^{(P_{4})}$YLv4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
$^{(P_{5})}$YLv5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
$^{(P_{6})}$YLv7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
$^{(P_{7})}$YLv10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
$^{(P_{8})}$YLv12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
$^{(P_{9})}$FRCNN(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
$^{(P_{10})}$DETR(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
$^{(P_{11})}$RT-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
$^{(P_{j})}$Source 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
@@ 4.1 Experimental Setup
We treat the person as the target object class, setting the batch size to 32, the maximum number of epochs $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, learning rates $\eta_{u}=0.01$ and $\eta_{p}=0.003$, and perturbation constraint $\epsilon=0.5$. We use the Adam optimizer [29] with its default parameters. Textual prompts $\mathcal{P}$ are automatically generated using BLIP [33]. The adopted version of Stable Diffusion [46] is v1.4. Since we generate irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$. To achieve stronger performance, we apply two iterations of optimization to the patch using BadPatch. We evaluate the effectiveness of different patch generation methods using both the Attack Success Rate (ASR) and Average Precision (AP). The ASR measures the success of the patches in misleading the detector, while AP reflects the overall decline in detection accuracy. The confidence threshold and IoU threshold for the detector are both set to 0.5. All experiments were conducted on a single NVIDIA A100 GPU.
@@ 4.2 Attack Performance Evaluation
We generate a series of adversarial patches targeting different models using BadPatch and compare them on the INRIA dataset with existing methods, including AdvYL [51], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34]. The detailed results are reported in Table 1. As shown, BadPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny. When compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YOLOv4 results in a 0.2% and 4.1% ASR increase, respectively. Additionally, as demonstrated in the Appendix B, BadPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
@@ 4.3 Cross-dataset Evaluation
We also evaluate an ensemble version of BadPatch, which trains the patch using averaged gradients from multiple models. We compare this with two other methods, NPAP and D2D, where NPAP was trained on YOLOv3 and YOLOv4, and D2D was trained on all considered models. Our ensemble BadPatch was similarly trained on all models. The results, presented in the Appendix C, demonstrate that our method outperforms NPAP and D2D by a significant margin.
@@ 4.4 Ablation Studies
To ensure the natural appearance of the generated patches, we constrain each optimization step using the $L_{\infty}$ norm. The choice of $\epsilon$ will influence visual naturalness of the generated patch and its attack performance. To quantify the natural appearance of the patches, we use GPT-4o [1] to generate descriptive text for reference images (denoted by ’-’) and compute the CLIP similarity [42] between these descriptions and the generated patches. The corresponding descriptions can be found in the Appendix D. As illustrated in Figure 4, we obtain patches with varying styles under different constraints using BadPatch and evaluate both the ASR and CLIP similarity for each. Based on the results, we select $\epsilon=0.5$ as the optimal value, as it achieves a favorable balance between attack performance and natural appearance.
@@
Here, we investigate the impact of patch shapes, as used for $P_{m}$ and $P_{n}$ in Table 1. We compare the performance of random noise images with square and irregular shapes (of the same size as BadPatch). From the ASR and AP results in Table 1, it is evident that square patches yield higher attack effectiveness compared to irregular patches. In other words, designing irregular patches with high attack success rates is more challenging.
@@ 4.5 Adversarial T-shirt Dataset
In total, we collected 1,131 images, including both individual and group photos (ranging from 2 to 10+ persons). We name the dataset AdvT-shirt-1K, which includes detailed annotations for person and patch locations (bounding boxes). The collection and annotation process of the dataset spanned six months, which involves the refinement of our BadPatch method for multiple times. The statistics of the dataset is shown in Figure 8 with several example images are provided in Figure 1. The results shown in the right subfigure of Figure 8 demonstrate the effectiveness of the adversarial T-shirts in evading the YOLOv5s detector. Although small in scale, our dataset provides a practical resource for developing and testing the real-world effectiveness of adversarial patch detectors.
@@ 5 Conclusion
In this work, we proposed a novel diffusion-based adversarial patch generation method BadPatch to generate naturalistic and customized patches based on a reference image. BadPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches. Extensive experiments across multiple object detection models validate the effectiveness of BadPatch. With BadPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of BadPatch in physical-world environments.